The flywheel of a steam engine of mass 1000 kg has radius of gyration as 1 m. If the maximum and minimum speed of the flywheel is 80 r.p.m. and 78 r.p.m. respectively, find the fluctuation of energy


[Ans, 1732.7 N-m).


1
Expert's answer
2021-02-18T18:39:54-0500
"E=0.5I(\omega_2^2-\omega_1^2)\\E=0.5(1000)(1)^2(80^2-78^2)\frac{4\pi^2}{60^2}\\E=1732.7\ J"